OpenStudy (anonymous):

The credit remaining on a phone card (in dollars) is a linear function of the total calling time made with the card (in minutes). The remaining credit after 42 minutes of calls is $14.12 , and the remaining credit after 74 minutes of calls is $9.64 . What is the remaining credit after 78 minutes of calls?

OpenStudy (anonymous):

ii got to $14.92-9.64=5.28 and 74-42=32 but now im lost

OpenStudy (anonymous):

it may be 9.64 - ( ($14.12 -9.64)/(74-42) *(78-74) )

OpenStudy (anonymous):

9.64-0.56=9.08

OpenStudy (anonymous):

what about this one @matricked Milan is driving to Boston. Suppose that the distance to his destination (in miles) is a linear function of his total driving time (in minutes). Milan has 68 miles to his destination after 19 minutes of driving, and he has 53.6 miles to his destination after 35 minutes of driving. How many miles will he have to his destination after 49 minutes of driving?

OpenStudy (anonymous):

it will be 53.6 - (((68-53.6) / (35-19)) * (49-35) )=53.6-12.6=41